### v1

# Geometric Property (T) and Kazhdan projections

Abstract.
We characterise Geometric Property (T) by the existence of a certain projection in the maximal uniform Roe algebra $C_{u,\max}^{*}(X)$, extending the notion of Kazhdan projection for groups to the realm of metric spaces. We also describe this projection in terms of the decomposition of the metric space into coarsely connected components.
Key words and phrases:
Geometric Property (T), Kazhdan projection, maximal uniform Roe algebra
2020 Mathematics Subject Classification:
Primary 22D55; Secondary 46L45, 51F30
This work is supported by the FONDECYT project 3230024.
1. Introduction
Geometric Property (T) was introduced in [6] as an analogue of Kazhdan’s Property (T) in the context of metric spaces. For groups, Property (T) can be characterised by the existence of a certain projection in the full group $*{}^{*}$-algebra; see [1, Lemma 2] or [4, Proposition 2]. In this paper, we extend this characterisation to metric spaces.
We will focus on metric spaces $(X,d)$ such that the distance $d$ is allowed to take the value $\infty$. We will always assume that $(X,d)$ has bounded geometry, meaning that, for every $R>0$,
$\sup_{x\in X}|B(x,R)|<\infty,$
where $B(x,R)$ denotes the ball of radius $R$ centred at $x$. A typical example of such a space is the disjoint union of a family of finite graphs of uniformly bounded degree. We define the tube of diameter $R$ as
$\operatorname{Tube}_{X}(R)=\{(x,y)\in X\times X\ \mid\ d(x,y)\leq R\}.$
We say that a subset $E$ of $X\times X$ is controlled if there is $R\in(0,\infty)$ such that $E\subseteq\operatorname{Tube}_{X}(R)$. Given two subsets $E,F\subset X\times X$, their composition is defined as
$E\circ F=\{(x,y)\in X\times X\ \mid\ \exists z\in X,\ (x,z)\in E,\ (z,y)\in F\}.$
We define inductively $E^{\circ(n+1)}=E\circ E^{\circ n}$, where $E^{\circ 1}=E$. We say that a controlled set $E$ is generating if, for any controlled set $F$, there is $n\geq 1$ such that $F\subseteq E^{\circ n}$. We say that $X$ is monogenic if there exists a controlled generating set inside $X$.
Two points $x,y$ are in the same coarse component if the set $\{(x,y)\}$ is controlled. This is an equivalence relation and the equivalence classes are called coarse components of $X$. The space is coarsely connected if there is only one coarse component. We will use the same convention as in [6]:
Convention: Throughout this paper, a space will mean a bounded geometry, monogenic metric space with at most countably many coarse components.
The translation algebra $\mathbb{C}_{u}[X]$ is the set of matrices $T=(T_{x,y})_{x,y\in X}$ with coefficients in $\mathbb{C}$, satisfying
$\sup_{x,y\in X}|T_{x,y}|<\infty,$
and such that the support
$\operatorname{supp}(T)=\{(x,y)\in X\times X\ \mid\ T_{x,y}\neq 0\}$
is a controlled set. The usual matrix operations and adjoint make $\mathbb{C}_{u}[X]$ into a $\ast$-algebra. The elements of $\mathbb{C}_{u}[X]$ are also called finite propagation operators.
A partial translation on $X$ is a bijective map $t:A\to B$, where $A,B$ are subsets of $X$, and such that the graph
$\operatorname{graph}(t)=\{(x,y)\in B\times A\ \mid\ x=t(y)\}$
is a controlled set. Partial translations can be viewed as elements of $\mathbb{C}_{u}[X]$ in the following way. If $t:A\to B$ is a partial translation, we can identify it with the matrix $v\in\mathbb{C}_{u}[X]$ defined by
$v_{x,y}=\begin{cases}1,&t(y)=x\\ 0,&\text{otherwise}.\end{cases}$ (1)
Observe that $\operatorname{supp}(v)=\operatorname{graph}(t)$. We will also refer to $v$ as a partial translation.
A representation of $\mathbb{C}_{u}[X]$ on a Hilbert space $\mathcal{H}$ is a unital $\ast$-homomorphism $\pi:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$, where $\mathcal{B}(\mathcal{H})$ stands for the algebra of bounded operators on $\mathcal{H}$. We say that $\xi\in\mathcal{H}$ is an invariant vector if
$\pi(v)\xi=\pi(vv^{*})\xi,$
for every partial translation $v$. We denote by $\mathcal{H}^{\pi}$ the subspace of invariant vectors of $\mathcal{H}$.
We say that a space $X$ has Property (T) if, for every controlled generating set $E\subseteq X$, there exists $c>0$ such that, for every representation $\pi:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$ and every $\xi\in(\mathcal{H}^{\pi})^{\perp}$, there is a partial translation $v$ with support in $E$ satisfying
$\|(\pi(v)-\pi(vv^{*}))\xi\|\geq c\|\xi\|.$
In order to characterise Property (T) in terms of projections, we need to consider the following subalgebra of $\mathbb{C}_{u}[X]$:
$\mathcal{A}_{X}=\left\{A\in\mathbb{C}_{u}[X]\,\mid\,\exists c\in\mathbb{C},\ \forall x\in X,\ \sum_{y\in X}A_{x,y}=\sum_{y\in X}A_{y,x}=c\right\}.$ (2)
This is a $\ast$-subalgebra of $\mathbb{C}_{u}[X]$, and the map $\theta_{X}:\mathcal{A}_{X}\to\mathbb{C}$ defined by
$\theta_{X}(A)=\sum_{y\in X}A_{x,y}$ (3)
is a $\ast$-homomorphism. We will denote by $\mathcal{A}_{X}^{+}$ the subset of $\mathcal{A}_{X}$ given by all those $A$ such that $A_{x,y}\geq 0$ for all $x,y\in X$. We also define $\left(\mathcal{A}_{X}^{+}\right)_{1}=\mathcal{A}_{X}^{+}\cap\theta_{X}^{-1}(\{1\})$.
Finally, the $*{}^{*}$-algebra $C_{u,\max}^{*}(X)$ is defined as the completion of $\mathbb{C}_{u}[X]$ for the norm
$\|T\|_{C_{u,\max}^{*}(X)}=\sup\left\{\|\pi(T)\|\ \mid\ \pi\text{ is a representation of }\mathbb{C}_{u}[X]\right\}.$
We call $C_{u,\max}^{*}(X)$ the maximal uniform Roe algebra of $X$. Recall that a projection in a $*{}^{*}$-algebra is an element $P$ such that $P=P^{*}=P^{2}$.
Theorem 1.1.
Let $X$ be a space. The following are equivalent:
  • i)
    The space $X$ has Property (T).
  • ii)
    There exists a projection $P$ in $C_{u,\max}^{*}(X)$ such that, for every representation $\pi:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$, $\pi(P)$ is the orthogonal projection onto the subspace of invariant vectors $\mathcal{H}^{\pi}$.
Moreover, in this case, $P$ lies in the closure of $\left(\mathcal{A}_{X}^{+}\right)_{1}$ inside $C_{u,\max}^{*}(X)$.
We see from the definition of the norm of $C_{u,\max}^{*}(X)$ that, if the projection in Theorem 1.1 exists, then it is unique. We call it the Kazhdan projection of $C_{u,\max}^{*}(X)$.
Since the spaces that we consider here decompose into coarsely connected components, it is natural to ask what the Kazhdan projection looks like on each component. More precisely, let
$X=\bigsqcup_{n}X_{n}$
be the decomposition of $X$ into coarsely connected components. Let $Q_{n}:\mathbb{C}_{u}[X]\to\mathbb{C}_{u}[X_{n}]$ be the restriction map to $X_{n}$. In other words,
$Q_{n}(T)=T\mathds{1}_{X_{n}},\quad\forall T\in\mathbb{C}_{u}[X],$
where $\mathds{1}_{X_{n}}$ is the partial translation associated to the identity on $X_{n}$. This map extends to a $\ast$-homomorphism $Q_{n}:C_{u,\max}^{*}(X)\to C_{u,\max}^{*}(X_{n})$; see Lemma 4.3.
Corollary 1.2.
Let $X$ be a space, and let $X=\bigsqcup_{n}X_{n}$ be its decomposition into coarsely connected components. If $X$ has Property (T), then so does $X_{n}$ for every $n$. Moreover, if $P$ denotes the Kazhdan projection of $C_{u,\max}^{*}(X)$, then $Q_{n}(P)$ belongs to $\mathbb{C}_{u}[X_{n}]$ and it is the Kazhdan projection of $C_{u,\max}^{*}(X_{n})$. Its coefficients are given by
[…]
for every $x,y$ in $X_{n}$.
This paper is organised as follows. In Section 2, we prove a particular case of the implication (ii)$\implies$(i) in Theorem 1.1. In Section 3, we prove the implication (i)$\implies$(ii), and finish the proof of Theorem 1.1. Corollary 1.2 is proved in Section 4.
Acknowledgements
This work originated from a series of very enlightening discussions that I had with Mikael de la Salle in Lyon during a research visit in February 2023. I am very grateful to him for his inspiring ideas, and to the members of the Institut Camille Jordan for their hospitality. I also thank Rufus Willett for his very valuable feedback.
2. A sufficient condition for Property (T)
We begin by showing that the existence of a Kazhdan projection in the closure of $\mathcal{A}_{X}$ implies Property (T). For this purpose, we need to review some facts about amenability.
2.1. Amenability
A space $X$ is said to be amenable if $\ell^{\infty}(X)$ admits an invariant mean. More precisely, there is a positive unital linear form $\phi:\ell^{\infty}(X)\to\mathbb{C}$ such that, for every $f\in\ell^{\infty}(X)$, and every partial translation $t:A\to B$ such that $B$ contains the support of $f$,
$\phi(f\circ t)=\phi(f).$
The following lemma will not be used in the proofs of our results. We state it here with the sole purpose of emphasising that, although this definition of amenability naturally extends the one for groups, it has some rather counter-intuitive properties.
Lemma 2.1.
Let $X$ be a space and let $X_{0}$ be a coarse component of $X$. If $X_{0}$ is amenable, then so is $X$.
Proof.
Let $\phi\in\ell^{\infty}(X_{0})^{*}$ be an invariant mean. Define $\tilde{\phi}:\ell^{\infty}(X)\to\mathbb{C}$ by
$\tilde{\phi}(f)=\phi(f|_{X_{0}}),\quad\forall f\in\ell^{\infty}(X),$
where $f|_{X_{0}}$ denotes the restriction of $f$ to $X_{0}$. Then $\tilde{\phi}$ is an invariant mean on $\ell^{\infty}(X)$. ∎
Observe that there is a copy of $\ell^{\infty}(X)$ inside $\mathbb{C}_{u}[X]$ given by the subalgebra of diagonal matrices. The following lemma is a reformulation of [6, Proposition 6.1]. Its proof relies on ultraproducts of spaces and operators; for details, we refer the reader to [3].
Lemma 2.2.
Let $X$ be a space. The following are equivalent:
  • (i)
    The space $X$ is amenable.
  • (ii)
    There exists a representation $\pi:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$ such that $\mathcal{H}^{\pi}\neq\{0\}$.
Proof.
Assume first that (ii) holds, and let $\xi$ be a unit vector in $\mathcal{H}^{\pi}$. We define a unital linear map $\phi:\mathbb{C}_{u}[X]\to\mathbb{C}$ by
$\phi(T)=\langle\pi(T)\xi,\xi\rangle,\quad\forall T\in\mathbb{C}_{u}[X].$
We claim that the restriction of $\phi$ to $\ell^{\infty}(X)$ is an invariant mean. Indeed, let $f\in\ell^{\infty}(X)$, $t:A\to B$ a partial translation such that $B$ contains the support of $f$, and let $v$ be the element of $\mathbb{C}_{u}[X]$ associated to $t$ as in (1). Observing that $f\circ t=v^{*}fv$ and $vv^{*}=\mathds{1}_{B}$, we see that
$\phi(f\circ t)$ $=\langle\pi(v^{*}fv)\xi,\xi\rangle$
$=\langle\pi(fv)\xi,\pi(v)\xi\rangle$
$=\langle\pi(f)\pi(vv^{*})\xi,\pi(vv^{*})\xi\rangle$
$=\langle\pi(f)\xi,\xi\rangle$
$=\phi(f).$
Finally, the identity $\phi(|f|^{2})=\|\pi(f)\xi\|^{2}$ shows that $\phi$ is positive. We conclude that $\phi$ is an invariant mean.
Conversely, assume that (i) holds. By [6, Lemma 6.1], there is a net of unit vectors $(\xi_{i})_{i\in I}$ in $\ell^{2}(X)$ such that, for every partial translation $v$,
$\|v\xi_{i}-vv^{*}\xi_{i}\|\to 0.$
Since $I$ is a directed set, there is a filter $\mathcal{F}$ on $I$ given by all the sets containing a subset of the form $\{\ i\in I\mid\ i\geq i_{0}\}$ for some $i_{0}\in I$. Let $\mathcal{U}$ be an ultrafilter containing $\mathcal{F}$, and let $\mathcal{H}=(\ell^{2}(X))_{\mathcal{U}}$ be the ultrapower Hilbert space. We can define a representation $\pi:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$ by
$\pi(T)(\eta_{i})_{\mathcal{U}}=(T\eta_{i})_{\mathcal{U}},\quad\forall T\in\mathbb{C}_{u}[X],\ \forall(\eta_{i})_{\mathcal{U}}\in\mathcal{H}.$
Then $\xi=(\xi_{i})_{\mathcal{U}}$ is a unit vector in $\mathcal{H}$, and
$\|\pi(v)\xi-\pi(vv^{*})\xi\|=\lim_{\mathcal{U}}\|v\xi_{i}-vv^{*}\xi_{i}\|=0.$
Therefore $\mathcal{H}^{\pi}$ is not trivial. ∎
In [6, Corollary 6.1], it was proved that, for infinite coarsely connected spaces, Property (T) is equivalent to non-amenability. Although the coarse connectedness assumption is needed for one direction of the equivalence, the other one holds in full generality. In order to prove this, we will use the following characterisation of Property (T); for a proof see [6, Proposition 3.1].
Lemma 2.3 (Willett–Yu).
Let $X$ be a space. The following are equivalent:
  • (i)
    The space $X$ has Property (T).
  • (ii)
    There exist a controlled generating set $E\subseteq X$ and $c>0$ such that, for every representation $\pi:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$ and every $\xi\in(\mathcal{H}^{\pi})^{\perp}$, there is a partial translation $v$ with support in $E$ satisfying
    $\|(\pi(v)-\pi(vv^{*}))\xi\|\geq c\|\xi\|.$
Remark 2.4.
Lemma 2.3 says that it suffices to consider one generating set in the definition of Property (T).
Lemma 2.5.
Let $X$ be a space. If $X$ does not have Property (T), then it is amenable.
Proof.
The proof follows the same ideas as that of Lemma 2.2. By Lemma 2.3, for every $n\geq 1$, there is a representation $\pi_{n}:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H}_{n})$ and a unit vector $\xi_{n}\in(\mathcal{H}_{n}^{\pi_{n}})^{\perp}$ such that, for every partial translation $v$ with support in $\operatorname{Tube}_{X}(n)$,
$\|(\pi_{n}(v)-\pi_{n}(vv^{*}))\xi_{n}\|<\frac{1}{n}.$
As in the proof of Lemma 2.2, taking an ultrafilter on $\mathbb{N}$ and considering the natural representation on the ultraproduct $(\mathcal{H}_{n})_{\mathcal{U}}$, we see that $\xi=(\xi_{n})_{\mathcal{U}}$ is a non-trivial invariant vector. By Lemma 2.2, $X$ is amenable. ∎
2.2. Proof of Property (T)
Now we will need a characterisation of Property (T) in terms of diagonal operators. Recall that there is a copy of $\ell^{\infty}(X)$ in $\mathbb{C}_{u}[X]$ given by the subalgebra of diagonal matrices. We can define a map $\Phi_{X}:\mathbb{C}_{u}[X]\to\ell^{\infty}(X)$ by
$\Phi_{X}(T)(x)=\sum_{y\in X}T_{x,y},\quad\forall T\in\mathbb{C}_{u}[X],\ \forall x\in X.$
The following was proved in [6, Lemma 3.1].
Lemma 2.6 (Willett–Yu).
Let $X$ be a space. For every representation $\pi:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$, a vector $\xi\in\mathcal{H}$ is invariant if and only if
$\pi(\Phi_{X}(T))\xi=\pi(T)\xi,$
for all $T\in\mathbb{C}_{u}[X]$.
Remark 2.7.
Recall the definition of $\mathcal{A}_{X}$ in (2). The image of $\mathcal{A}_{X}$ under $\Phi_{X}$ consists of the constant functions on $X$. Moreover, for all $A\in\mathcal{A}_{X}$ and $x\in X$,
$\Phi_{X}(A)(x)=\theta_{X}(A),$ (4)
where $\theta_{X}$ is the morphism defined in (3).
The following characterisation is contained in [6, Proposition 3.1].
Proposition 2.8 (Willett–Yu).
Let $X$ be a space. The following are equivalent:
  • (i)
    The space $X$ does not have Property (T).
  • (ii)
    For every controlled generating set $E$ and every $\varepsilon>0$, there is a representation $\pi:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$ and $\xi\in(\mathcal{H}^{\pi})^{\perp}$ with $\|\xi\|=1$ such that, for every $T\in\mathbb{C}_{u}[X]$ with support in $E$, we have
    $\left\|\pi(T)\xi-\pi(\Phi_{X}(T))\xi\right\|\leq\varepsilon\sup_{x,y\in X}|T_{x,y}|.$
Now we can prove a particular case of the implication (ii)$\implies$(i) in Theorem 1.1.
Lemma 2.9.
Let $X$ be a space such that there is a projection $P$ in the closure of $\mathcal{A}_{X}$ inside $C_{u,\max}^{*}(X)$ such that, for every representation $\pi:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$, $\pi(P)$ is the orthogonal projection onto the subspace of invariant vectors $\mathcal{H}^{\pi}$. Then $X$ has Property (T).
Proof.
First observe that $P$ is 00 if and only if $\mathbb{C}_{u}[X]$ does not admit a representation with non-trivial invariant vectors. By Lemma 2.2, this is equivalent to non-amenability, which in turn implies Property (T) by Lemma 2.5. Assume now that $P\neq 0$, and let $(A_{n})$ be a sequence in $\mathcal{A}_{X}$ such that
$\|A_{n}-P\|_{C_{u,\max}^{*}(X)}\to 0.$
Let $\rho:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$ be a representation such that $\mathcal{H}^{\rho}$ is not trivial. Take $\eta\in\mathcal{H}^{\rho}$ with $\|\eta\|=1$, and recall that $\rho(A_{n})\eta=\rho(\Phi_{X}(A_{n}))\eta$ by Lemma 2.6. Observing that $\rho(\Phi_{X}(A_{n}))=\theta_{X}(A_{n})\operatorname{Id}_{\mathcal{H}}$, we see that
$\theta_{X}(A_{n})$ $=\theta_{X}(A_{n})\langle\eta,\eta\rangle$
$=\langle\rho(\Phi_{X}(A_{n}))\eta,\eta\rangle$
$=\langle\rho(A_{n})\eta,\eta\rangle$
$n\to\infty]{}\langle\rho(P)\eta,\eta\rangle$
$=1.$
Normalising this sequence, we find $A$ in $\mathcal{A}_{X}$ such that $\theta_{X}(A)=1$ and
$\|A-P\|_{C_{u,\max}^{*}(X)}\leq\frac{1}{3}.$
Assume by contradiction that $X$ does not have Property (T), and define
$C=\sup_{x,y\in X}|A_{x,y}|.$
By Proposition 2.8, there is a representation $\pi:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$ and $\xi\in(\mathcal{H}^{\pi})^{\perp}$ with $\|\xi\|=1$ such that
$\left\|\pi(A)\xi-\pi(\Phi_{X}(A))\xi\right\|\leq\frac{1}{3C}\sup_{x,y\in X}|A_{x,y}|=\frac{1}{3}.$
Since $\theta_{X}(A)=1$, we have $\pi(\Phi_{X}(A))\xi=\xi$ and therefore
$\left\|\pi(A)\xi-\xi\right\|\leq\frac{1}{3}.$
Finally, since $\pi(P)\xi=0$, we obtain
$\|\xi\|$ $=\|\xi-\pi(P)\xi\|$
$\leq\|\xi-\pi(A)\xi\|+\|\pi(A)\xi-\pi(P)\xi\|$
$\leq\frac{2}{3},$
which is a contradiction. ∎
3. Kazhdan projections and Property (T)
In this section, we finish the proof of Theorem 1.1. For the proof of the implication (i)$\implies$(ii), we will need some facts about colourings on graphs.
3.1. Graphs and colourings
Let $X=(V,E)$ be a graph and $d\geq 1$ an integer. We say that a map $c:E\to\{1,\ldots,d\}$ is a $d$-colouring of $E$ if, whenever $e_{1},e_{2}$ are two adjacent edges, we have $c(e_{1})\neq c(e_{2})$. The following result was proved in [5]; see also [2, Theorem 5.3.2].
Theorem 3.1 (Vizing).
Let $d\in\mathbb{N}$ and let $X=(V,E)$ be a finite graph such that every vertex in $X$ has degree at most $d$. Then there exists a $(d+1)$-colouring of $E$.
This result extends to infinite graphs by standard arguments. We include the proof of this fact for completeness.
Corollary 3.2.
Let $d\in\mathbb{N}$ and let $X=(V,E)$ be a countable infinite graph such that every vertex in $X$ has degree at most $d$. Then there exists a $(d+1)$-colouring of $E$.
Proof.
Let $(X_{n})$ be an increasing sequence of finite subgraphs of $X$ such that $X=\bigcup_{n}X_{n}$. Let us write $X_{n}=(V_{n},E_{n})$. Since $E_{n}$ is contained in $E$, every $v\in V_{n}$ has degree at most $d$ in $X_{n}$. By Theorem 3.1, $E_{n}$ admits a $(d+1)$-colouring. Let
$K=\{1,\ldots,d+1\}^{E}.$
Recall that this is a compact space for the product topology. For each $n$, define
[…]
Then $(K_{n})$ is a decreasing sequence of closed subsets of $K$. Moreover, they are non-empty by the observation above. By the finite intersection property, the set $\bigcap_{n}K_{n}$ is non-empty. Any element of this intersection is a $(d+1)$-colouring of $E$. ∎
Let now $X$ be a space. We denote by $\mathcal{S}_{X}$ the set of finite propagation permutation matrices. In other words, every $A$ in $\mathcal{S}_{X}$ is an element of $(\mathcal{A}_{X}^{+})_{1}$ with coefficients in $\{0,1\}$.
Lemma 3.3.
Let $X$ be a space and let $R>0$. There exist $A_{0},A_{1},\ldots,A_{n}\in\mathcal{S}_{X}$ such that, for every partial translation $v$ with support in $\operatorname{Tube}_{X}(R)$, there exist $f_{0},f_{1},\ldots,f_{n}\in\ell_{\infty}(X)$ with $f_{i}^{2}=f_{i}$ such that
$v=\sum_{i=0}^{n}f_{i}A_{i}.$
Moreover, we have $vv^{*}=\sum_{i=0}^{n}f_{i}$.
Proof.
Define a graph $(V,E)$ with $V=X$ and
$E=\left\{\{x,y\}\ \mid\ x\neq y,\ d(x,y)\leq R\right\}.$
It has bounded degree because $X$ has bounded geometry. By Corollary 3.2, there is a colouring $c:E\to\{1,\ldots,n\}$. We will extend $c$ to the diagonal of $X\times X$ by
$c(\{x,x\})=0,\quad\forall x\in X.$
Define $A_{0}=1$ and $A_{i}$ ($i=1,\ldots,n$) by
$(A_{i})_{x,y}=\begin{cases}1&\text{if }\{x,y\}\in c^{-1}(\{i\}),\\ 1&\text{if }x=y,\ \nexists\,z\text{ such that }\{x,z\}\in c^{-1}(\{i\}),\\ 0&\text{otherwise}.\end{cases}$
Now, given a partial translation $v$ with support in $\operatorname{Tube}_{X}(R)$, we define $f_{i}$ ($i=0,\ldots,n$) by
$f_{i}(x)=\begin{cases}1&\text{if }\exists\,y\text{ such that }\{x,y\}\in c^{-1}(\{i\})\text{ and }v_{x,y}=1,\\ 0&\text{otherwise}.\end{cases}$
Then we have, for every $i\in\{0,\ldots,n\}$,
$f_{i}(x)(A_{i})_{x,y}=\begin{cases}1&\text{if }\{x,y\}\in c^{-1}(\{i\})\text{ and }v_{x,y}=1,\\ 0&\text{otherwise},\end{cases}$
which shows that
$v=\sum_{i=0}^{n}f_{i}A_{i}.$
Moreover,
$\sum_{i=0}^{n}f_{i}(x)=\begin{cases}1&\text{if }\exists y,\ v_{x,y}=1,\\ 0&\text{otherwise},\end{cases}$
which means that $vv^{*}=\sum_{i=0}^{n}f_{i}$. ∎
3.2. Proof of the necessary condition
Now we prove the implication (i)$\implies$(ii) in Theorem 1.1.
Lemma 3.4.
Let $X$ be a space with Property (T). Then there exists a projection $P$ in the closure of $(\mathcal{A}_{X}^{+})_{1}$ inside $C_{u,\max}^{*}(X)$ such that, for every representation $\pi:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$, $\pi(P)$ is the orthogonal projection onto the subspace of invariant vectors $\mathcal{H}^{\pi}$.
Proof.
Let $E$ be a controlled generating set. By definition, there is a constant $c>0$ such that, for every representation $\pi:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$ and every $\xi\in(\mathcal{H}^{\pi})^{\perp}$, there is a partial translation $v$ with support in $E$ such that
$\|\pi(v)\xi-\pi(vv^{*})\xi\|\geq c\|\xi\|.$ (5)
By Lemma 3.3, there exist $A_{0},A_{1},\ldots,A_{n}\in\mathcal{S}_{X}$ such that, for every such $v$, there exist $f_{0},f_{1},\ldots,f_{n}\in\ell_{\infty}(X)$ satisfying $f_{i}^{2}=f_{i}$ and
$v=\sum_{i=0}^{n}f_{i}A_{i}.$
Since $\pi(vv^{*})=\sum\pi(f_{i})$ and $\|\pi(f_{i})\|^{2}=\|\pi(f_{i})\|$, the inequality (5) yields
$c\|\xi\|$ $\leq\left\|\sum_{i=1}^{n}\pi(f_{i})(\pi(A_{i})\xi-\xi)\right\|$
$\leq\sum_{i=1}^{n}\|\pi(A_{i})\xi-\xi\|.$
This shows that there exists $j$ in $\{1,\ldots,n\}$ such that
$\|\pi(A_{j})\xi-\xi\|\geq\frac{c}{n}\|\xi\|.$
Observe that $\pi(A_{j})$ is a unitary operator because $A_{j}$ is a unitary matrix in $\mathbb{C}_{u}[X]$. Hence, by the parallelogram identity,
$\left\|\frac{\pi(A_{j})\xi+\xi}{2}\right\|\leq\delta\|\xi\|,$
where $\delta=\left(1-\left(\frac{c}{2n}\right)^{2}\right)^{1/2}$. Notice that $\delta$ only depends on the controlled set $E$, and not on the representation $\pi$ or the vector $\xi$. Defining $A\in\left(\mathcal{A}_{X}^{+}\right)_{1}$ by
$A=\frac{1}{n}\sum_{i=1}^{n}\frac{1+A_{i}}{2},$
we see that
$\|\pi(A)\xi\|$ $\leq\frac{1}{n}\left\|\frac{\xi+\pi(A_{j})\xi}{2}\right\|+\frac{1}{n}\sum_{i\neq j}\left\|\frac{\xi+\pi(A_{i})\xi}{2}\right\|$
$\leq\frac{\delta}{n}\|\xi\|+\frac{n-1}{n}\|\xi\|$
$=\tilde{\delta}\|\xi\|,$
where $\tilde{\delta}=1-\frac{1-\delta}{n}<1$. This inequality holds for every representation $\pi$ and every $\xi$ in $(\mathcal{H}^{\pi})^{\perp}$. Now let $p_{\pi}$ be the orthogonal projection onto $\mathcal{H}^{\pi}$. Observe that $\pi(A)p_{\pi}=p_{\pi}$ because $\theta_{X}(A)=1$. On the other hand,
$p_{\pi}\pi(A)=\left(\pi(A^{*})p_{\pi}\right)^{*}=p_{\pi}^{*}=p_{\pi}.$
Using these two facts, an induction argument shows that
$\pi(A^{k})-p_{\pi}=(\pi(A)-p_{\pi})^{k},$
for all $k\geq 1$. Let $\eta\in\mathcal{H}$. By the discussion above, we know that
$\|\pi(A)\eta-p_{\pi}\eta\|$ $=\|\pi(A)(\eta-p_{\pi}\eta)\|$
$\leq\tilde{\delta}\|\eta-p_{\pi}\eta\|$
$\leq\tilde{\delta}\|\eta\|.$
In other words, $\|\pi(A)-p_{\pi}\|\leq\tilde{\delta}$. Furthermore, for all $k\geq 1$,
$\left\|\pi(A^{k})-p_{\pi}\right\|=\left\|(\pi(A)-p_{\pi})^{k}\right\|\leq\tilde{\delta}^{k}.$ (6)
Since this bound holds for every representation, we see that $(A^{k})_{k\geq 1}$ is a Cauchy sequence in $C_{u,\max}^{*}(X)$, and therefore it converges to an element $P$. Since $A^{k}$ belongs to $\left(\mathcal{A}_{X}^{+}\right)_{1}$ for every $k\geq 1$, $P$ is in the closure of $\left(\mathcal{A}_{X}^{+}\right)_{1}$. The inequality (6) shows that $\pi(P)=p_{\pi}$ for every representation $\pi$, which in turn shows that $P$ is itself a projection. ∎
Observe that Lemmas 2.9 and 3.4 yield a particular case of Theorem 1.1. More precisely, a space $X$ has Property (T) if and only if there is a Kazhdan projection in the closure of $\mathcal{A}_{X}$ inside $C_{u,\max}^{*}(X)$.
3.3. Proof of the sufficient condition
Now we will show that, if $C_{u,\max}^{*}(X)$ has a Kazhdan projection, then it necessarily lies in the closure of $\mathcal{A}_{X}$, completing the proof of Theorem 1.1.
Lemma 3.5.
Let $X$ be a space, and assume that there is $P$ in $C_{u,\max}^{*}(X)$ such that, for every representation $\pi:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$, $\pi(P)$ is the orthogonal projection onto the subspace of invariant vectors $\mathcal{H}^{\pi}$. Then $X$ has Property (T).
Proof.
If $P$ belongs to the closure of $\mathcal{A}_{X}$, then we know that $X$ has Property (T) thanks to Lemma 2.9. Assume now that $P$ belongs to $C_{u,\max}^{*}(X)\setminus\overline{\mathcal{A}_{X}}$. By Lemma 3.4, $X$ does not have Property (T) since otherwise $P$ would belong to the closure of $(\mathcal{A}_{X}^{+})_{1}$. Thus, by Lemma 2.3, for every $n\geq 1$, there is a representation $\pi_{n}:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H}_{n})$ and a unit vector $\xi_{n}\in(\mathcal{H}_{n}^{\pi_{n}})^{\perp}$ such that, for every partial translation $v$ with support in $\operatorname{Tube}_{X}(n)$,
$\|(\pi_{n}(v)-\pi_{n}(vv^{*}))\xi_{n}\|<\frac{1}{n}.$
As in the proof of Lemma 2.2, taking an ultrafilter on $\mathbb{N}$ and considering the natural representation on the ultraproduct $(\mathcal{H}_{n})_{\mathcal{U}}$, we see that $\xi=(\xi_{n})_{\mathcal{U}}$ is a non-trivial invariant vector. However, since $\xi_{n}$ belongs to $(\mathcal{H}_{n}^{\pi_{n}})^{\perp}$, we have
$1=\|\pi(P)\xi\|=\lim_{\mathcal{U}}\|\pi_{n}(P)\xi_{n}\|=0,$
which is a contradiction. Therefore $P$ belongs to $\overline{\mathcal{A}_{X}}$, and $X$ has Property (T) by Lemma 2.9. ∎
Proof of Theorem 1.1.
The proof follows directly from Lemmas 3.4 and 3.5. ∎
4. Kazhdan projections and coarsely connected components
In this section, we study coarsely connected spaces and prove Corollary 1.2. We begin by characterising the Kazhdan projection of a finite space.
Lemma 4.1.
Let $X$ be a finite, coarsely connected space. Then $X$ has Property (T), and its Kazhdan projection $P$ is given by
$P_{x,y}=\frac{1}{|X|},\quad\forall x,y\in X.$
Proof.
Let $P\in\mathbb{C}_{u}[X]$ be given by $P_{x,y}=|X|^{-1}$ for all $x,y\in X$. Observe that $P$ belongs to $(\mathcal{A}_{X}^{+})_{1}$, and $P=P^{*}=P^{2}$. By Theorem 1.1, we only need to show that $\pi(P)$ is the projection onto $\mathcal{H}^{\pi}$ for every representation $\pi:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$. First, if $\xi$ belongs to $\mathcal{H}^{\pi}$, then
$\pi(P)\xi=\pi(\Phi_{X}(P))\xi=\pi(1)\xi=\xi.$
On the other hand, for all $T\in\mathbb{C}_{u}[X]$ and $x,y\in X$,
$(TP)_{x,y}=\sum_{z\in X}T_{x,z}|X|^{-1}=\Phi_{X}(T)_{x,x}|X|^{-1}=\sum_{z\in X}\Phi_{X}(T)_{x,z}P_{z,y}=(\Phi_{X}(T)P)_{x,y}.$
Hence, for every $\xi\in\mathcal{H}$,
$\pi(T)\pi(P)\xi=\pi(\Phi_{X}(T))\pi(P)\xi,$
which shows that $\pi(P)\xi$ is an invariant vector. We conclude that $\pi(P)$ is the orthogonal projection onto $\mathcal{H}^{\pi}$. ∎
Lemma 4.1 settles the finite case. The infinite case is even simpler.
Lemma 4.2.
Let $X$ be an infinite, coarsely connected space. If $X$ has Property (T), then it is not amenable. In particular, the Kazhdan projection of $C_{u,\max}^{*}(X)$ is 00.
Proof.
Assume by contradiction that $X$ is amenable. By [6, Lemma 6.1], there is a net of unit vectors $(\xi_{i})_{i\in I}$ in $\ell^{2}(X)$ such that, for every partial translation $v$,
$\|v\xi_{i}-vv^{*}\xi_{i}\|\to 0.$
Let $\lambda:C_{u,\max}^{*}(X)\to\mathcal{B}(\ell^{2}(X))$ be the extension of the natural representation of $\mathbb{C}_{u}[X]$ on $\ell^{2}(X)$. As in the proof of Lemma 2.2, we take an ultrafilter $\mathcal{U}$ on $I$, and consider the representation
$\pi(T)(\eta_{i})_{\mathcal{U}}=(\lambda(T)\eta_{i})_{\mathcal{U}},\quad\forall T\in C_{u,\max}^{*}(X),\ \forall(\eta_{i})_{\mathcal{U}}\in\mathcal{H},$
on the space $\mathcal{H}=(\ell^{2}(X))_{\mathcal{U}}$. Then $\xi=(\xi_{i})_{\mathcal{U}}$ is a non-trivial invariant vector for $\pi$. In other words, if $P$ denotes the Kazhdan projection of $C_{u,\max}^{*}(X)$, then $\pi(P)\xi=\xi\neq 0$. On the other hand, since $X$ is infinite and coarsely connected, $\ell^{2}(X)$ does not have non-trivial invariant vectors. In particular, $\lambda(P)\xi_{i}=0$. This shows that
$\|\xi\|=\|\pi(P)\xi\|=\lim_{\mathcal{U}}\|\lambda(P)\xi_{i}\|=0,$
which is a contradiction. We conclude that $X$ is not amenable. Therefore, by Lemma 2.2, $P$ is 00. ∎
Now we have completely characterised Kazhdan projections of coarsely connected spaces. In order to prove Corollary 1.2, we need to study the restrictions of a Kazhdan projection on a space to its coarse components.
Lemma 4.3.
Let $X$ be a space, and let $X_{0}$ be a coarse component of $X$. The map $Q:\mathbb{C}_{u}[X]\to\mathbb{C}_{u}[X_{0}]$ defined by
$T\mapsto T\mathds{1}_{X_{0}},\quad\forall T\in\mathbb{C}_{u}[X],$
where $\mathds{1}_{X_{0}}$ is the indicator function of $X_{0}$ seen as a diagonal operator, extends to a unital normal-$\ast$-homomorphism $Q:C_{u,\max}^{*}(X)\to C_{u,\max}^{*}(X_{0})$.
Proof.
Observe that $Q$ is a unital $\ast$-homomorphism from $\mathbb{C}_{u}[X]$ to $\mathbb{C}_{u}[X_{0}]$. We only need to show that it extends to a continuous map from $C_{u,\max}^{*}(X)$ to $C_{u,\max}^{*}(X_{0})$. Let $T\in\mathbb{C}_{u}[X]$, and let us write it by blocks as follows:
$T=T_{0}\oplus T^{\prime},$
where $T_{0}\in\mathbb{C}_{u}[X_{0}]$ and $T^{\prime}\in\mathbb{C}_{u}[X\setminus X_{0}]$. Observe that $Q(T)=T_{0}$. Let us now take a representation $\pi:\mathbb{C}_{u}[X_{0}]\to\mathcal{B}(\mathcal{H})$, and let $\tilde{\pi}:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H}\oplus\ell^{2}(X\setminus X_{0}))$ be given by
$\tilde{\pi}(T_{0}\oplus T^{\prime})=\pi(T_{0})\oplus T^{\prime}.$
Then
$\|\pi(T_{0})\|_{\mathcal{B}(\mathcal{H})}$ $\leq\max\left\{\|\pi(T_{0})\|_{\mathcal{B}(\mathcal{H})},\|T^{\prime}\|_{\mathcal{B}(\ell^{2}(X\setminus X_{0}))}\right\}$
$=\|\tilde{\pi}(T_{0}\oplus T^{\prime})\|_{\mathcal{B}(\mathcal{H}\oplus\ell^{2}(X\setminus X_{0}))}$
$\leq\|T_{0}\oplus T^{\prime}\|_{C_{u,\max}^{*}(X)}.$
Taking the supremum over all the representations of $\mathbb{C}_{u}[X_{0}]$, we find
$\|Q(T)\|_{C_{u,\max}^{*}(X_{0})}\leq\|T\|_{C_{u,\max}^{*}(X)},$
which completes the proof. ∎
Lemma 4.4.
Let $X$ be a space, $X_{0}$ a coarse component of $X$, and $Q$ the map defined in Lemma 4.3. If $X$ has Property (T), then so does $X_{0}$. Moreover, if $P$ is the Kazhdan projection of $C_{u,\max}^{*}(X)$, then $Q(P)$ is the Kazhdan projection of $C_{u,\max}^{*}(X_{0})$.
Proof.
Since $Q$ is a $\ast$-homomorphism, $Q(P)$ is a projection. Now let $\pi:\mathbb{C}_{u}[X_{0}]\to\mathcal{B}(\mathcal{H})$ be a representation, and define $\tilde{\pi}:\mathbb{C}_{u}[X]\to\mathcal{B}(\mathcal{H})$ by $\tilde{\pi}=\pi\circ Q$. Again, since $Q$ is a $\ast$-homomorphism, $\tilde{\pi}$ is indeed a representation. Furthermore, $\tilde{\pi}(P)$ is the orthogonal projection onto the subspace of invariant vectors $\mathcal{H}^{\tilde{\pi}}$. Since $\tilde{\pi}(P)=\pi(Q(P))$, we only need to show that $\mathcal{H}^{\tilde{\pi}}=\mathcal{H}^{\pi}$. This is true because
$\xi\in\mathcal{H}^{\tilde{\pi}}$ $\iff\tilde{\pi}(T)\xi=\tilde{\pi}(\Phi_{X}(T))\xi,\quad\forall T\in\mathbb{C}_{u}[X]$
$\iff\pi(Q(T))\xi=\pi(Q(\Phi_{X}(T)))\xi,\quad\forall T\in\mathbb{C}_{u}[X]$
$\iff\pi(T_{0})\xi=\pi(\Phi_{X_{0}}(T_{0}))\xi,\quad\forall T_{0}\in\mathbb{C}_{u}[X_{0}]$
$\iff\xi\in\mathcal{H}^{\pi}.$
Now we are ready to prove Corollary 1.2.
Proof of Corollary 1.2.
The fact that each coarse component $X_{n}$ has Property (T) follows from Lemma 4.4. Moreover, Lemma 4.4 also implies that $Q_{n}(P)$ is the Kazhdan projection of $C_{u,\max}^{*}(X_{n})$. The description of $Q_{n}(P)$ in each case follows from Lemmas 4.1 and 4.2. ∎